OpenStudy (anonymous):

Indefinite integral of: sqrt(1-x^2)dx Without trig substitution..

OpenStudy (anonymous):

Try multiplying integrand by \[\frac{ \sqrt{1-x^2} }{ \sqrt{1-x^2} }\] then use integration by parts